OpenStudy (anonymous):

when solving ft = (m1vf - m2vi) is the answer kgm/s or nxs (n times s)

OpenStudy (surry99):

Hint: perhaps the units are equivalent

OpenStudy (anonymous):

Just look at the units of equation. m1 and m2 are masses, so kg, Vi and Vf are velocities, so m/s. The answer is what you get when you multiply these together, and they are multiplied together in the equation. m1*vf = kg * m/s = kgm/s.

OpenStudy (anonymous):

it might also help to remember that n*s is equivalent to kgm/s.